OpenStudy (anonymous):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n2 + 2,500n - 5,250 Part A: What are the zeroes of the above function, and what do they represent? Show your work. (4 points) Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. (4 points) Part C: What is the axis of symmetry of the function P(n)? (2 points

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@$w3G_Godd

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

A) P(n) means the profit of selling tickets at n price. So, say P(12) = 10,000. This means the profit is $10,000 if the they decide to sell the concert tickets at $12 each. find zeros means find what value of n gives P(n) = 0. suppose n is, like P(500) = 0, this means there would be no profit if each ticket cost like $500 each. Of course you to find the actual value of n that makes the equations equal 0.

OpenStudy (anonymous):

B and C are just calculator problems. ill leave that to @ganeshie8

OpenStudy (anonymous):

Wait so I'm not understanding what the zeros mean?

OpenStudy (anonymous):

@$w3G_Godd

OpenStudy (anonymous):

maybe @ganeshie8 can explain it better for me im not really good at that

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I think i got b and c. I just need a

OpenStudy (anonymous):

for b i gotP(n)=-250n^2 +2,500n-4,000 p(n)'= -500n + 2500 pur p(n)'=0 u will get n = 5. -250(5^2) + 2500(5) - 4000 = 2250

OpenStudy (anonymous):

for c i got(b/2*a) b=2500 and a=-250 -(b/2*a)=(2500/500) =5

ganeshie8 (ganeshie8):

check your messages

OpenStudy (anonymous):

thank you so much!